# Traffic Scenario Logic: A Spatial-Temporal Logic for Modeling and Reasoning of Urban Traffic Scenarios

### v1

@@ Abstract
Formal representations of traffic scenarios can be used to generate test cases for the safety verification of autonomous driving. However, most existing methods are limited in highway or highly simplified intersection scenarios due to the intricacy and diversity of traffic scenarios. In response, we propose Traffic Scenario Logic (TSL), which is a spatial-temporal logic designed for modeling and reasoning of urban pedestrian-free traffic scenarios. TSL provides a formal representation of the urban road network that can be derived from OpenDRIVE, i.e., the de facto industry standard of high-definition maps for autonomous driving, enabling the representation of a broad range of traffic scenarios. We implemented the reasoning of TSL using Telingo, i.e., a solver for temporal programs based on the Answer Set Programming, and tested it on different urban road layouts. Demonstrations show the effectiveness of TSL in test scenario generation and its potential value in areas like decision-making and control verification of autonomous driving.
@@ 1 Introduction
Autonomous driving technology is promising to revolutionize transportation by enhancing safety, efficiency, and accessibility. The rigorous testing and verifying their safety under various traffic scenarios is central to developing and deploying autonomous vehicles. Formal representations of these scenarios are crucial for generating comprehensive test cases and ensuring the reliability of autonomous driving systems. (Zhang et al., 2021)
Existing methods for formalizing traffic scenarios often face limitations, particularly in modeling intricate urban environments with diverse traffic patterns. While some approaches excel in representing highway scenarios or simplified intersections, they struggle to capture the complexity inherent in urban driving environments. (Hilscher and Schwammberger, 2016; Yao and Li, 2023; Li, Tao, and Wotawa, 2020) Consequently, there is a pressing need for a versatile and robust formal logic framework capable of modeling a broad spectrum of urban traffic scenarios.
In response to this challenge, we introduce Traffic Scenario Logic (TSL), a novel spatial-temporal logic specifically tailored for modeling and reasoning about urban pedestrian-free traffic scenarios. Unlike previous methods, TSL is designed to provide a comprehensive formal representation of the urban road network, enabling the modeling of various traffic configurations encountered in complex urban environments.
One of the main benefits of TSL is its compatibility with OpenDRIVE, which is the industry-standard format used for high-definition (HD) maps in autonomous driving. (Dupuis, Strobl, and Grezlikowski, 2010) By utilizing OpenDRIVE data, TSL seamlessly translates real-world road networks into logical representations, enabling the representation of a wide range of urban traffic scenarios.
TSL also offers a spatial representation that sets it apart from previous methods. Unlike other methods that use a simple discretized position description, TSL represents the longitudinal position of a vehicle by its positional relationship with other cars. This is inspired by homotopy classes in trajectory planning, where a homotopy class is the set of all possible trajectories from a start to a goal point that avoids obstacles in the same manner. Similarly, we can categorize two scenarios into the same homotopy class if the positional relationships of vehicles in them vary in the same way. Distinguishing between scenarios of different homotopy classes ensures comprehensiveness, while merging same-class scenarios improves efficiency.
We first give the syntax and semantics of TSL for uni-direction multi-lane highway scenarios in Section 4, then extend it to urban scenarios in Section 5. Finally, we implement the reasoning of TSL with Telingo (Cabalar et al., 2019), a solver for temporal programs based on the Answer Set Programming. Section 6 demonstrates several examples on different urban road layouts, showing the effectiveness of TSL in test scenario generation and potential applications in decision-making and control verification of autonomous driving.
@@
  • Provides a formal representation of an urban road network that is compatible with the OpenDRIVE HD map format and a spatial representation that corresponds to homotopy classes of scenarios.
  • Implements the reasoning mechanism with Telingo solver and provides application examples on different road layouts.
@@ 2 Related Work
Hilscher et al. (2011) were one of the first few work on using spatial-temporal logic to model traffic scenarios, proposing Multi-Lane Spatial Logic (MLSL) which provides an abstract model of a uni-directional multi-lane motorway. In MLSL, vehicles’ position, speed, and acceleration constitute a traffic snapshot, and the traffic snapshots at different moments constitute a traffic scenario. Changes between traffic snapshots are described by transition rules, including acceleration and lane changes made by the vehicle, as well as changes in speed and position over time, inspired by $\mathcal{ITL}$ (Moszkowski, 1982). Hilscher, Linker, and Olderog (2013) proposed an extended version of MLSL that supports a bi-directional motorway and length measurement for road segments. Hilscher and Schwammberger (2016) further proposed Urban Multi-Lane Spatial Logic (UMLSL) which contains the modeling of intersections. UMLSL simplifies an intersection of two two-way, single-lane highways into four crossing road segments. Meanwhile, Xu and Li (2016) introduced vertical lanes that intersect with the horizontal lanes by extending the position to two dimensions to create an urban road network, formalized in a grid way. Based on UMLSL, Schwammberger and Alves (2021) proposed Urban Spatial Logic for Traffic Rules (USL-TR) to add static objects and non-autonomous road users. Xu et al. (2019) represented an urban road network with road segments and the arcs that link them together, which was further formalized into a hierarchical structure by Yao and Li (2023). This structure can express complex road networks like roundabouts. Some other research includes modeling an intersection controlled by traffic lights (Loos and Platzer, 2011); using the State-Clock Logic for the representation of time (Bischopink and Olderog, 2022); and employing ontology (Bagschik, Menzel, and Maurer, 2018; Li, Tao, and Wotawa, 2020).
@@
The above methods for describing spatial relationships rely on either road segments or a grid-based occupancy approach, which have their limitations. In most cases, distinguishing between consecutive multiple free grids or road segments is not necessary. On the other hand, there are instances when it is crucial to describe the positional relationships of vehicles within the same grid or road segment. In general, a simple discretized description of the road network is insufficient to meet the needs of test scenario generation.
3 Preliminary
In this section, we provide a brief introduction to the road coordinate system and OpenDRIVE map format, the de facto industry standard of HD maps for autonomous driving.
The road coordinate system, also known as the Frenet frame, describes motions on a continuous differentiable curve in 3-dimensional Euclidean space. Since the slope angle and superelevation of a road in the vertical direction are commonly small, what is used for road coordinate description is mostly a 2-dimensional version. The road coordinate system is composed of the $S$ axis along the direction of the road reference line (generally the road centerline) and its orthogonal $D$ axis (also referred to as the $L$ or $T$ axis in some literature). The coordinates of a point on the road are determined by its accumulated distance $s$ along the road reference line from the start point and the distance $d$ between the point and its projection onto the reference line. With the road coordinate system, we can mitigate the influence of road curvature and uniformly describe the vehicle’s position on the road.
reference line $d$$s$$S$$D$projection point
Figure 1: The road coordinate system. The coordinate of the red dot consists of its accumulated distance $s$ along the reference line and the distance $d$ to the projection point.
OpenDRIVE [^1] https://www.asam.net/standards/detail/opendrive/ is a standardized HD map file format for representing road networks, encapsulating information about road geometry, topology, traffic signs, etc. Each road in an OpenDRIVE file consists of road segments with a specific number of lanes, as depicted in 2(a). Junctions are used to connect the roads, creating a connecting road for each pair of connected lanes. Due to its powerful representation capabilities, OpenDRIVE has become the most widely used HD map format in autonomous driving and simulators such as CARLA.
Junction
(a) OpenDRIVE map
connect pointintersect pointl1l2l3l4l5l6l13l42l45l62l63l15n1n2n3n4n5n6x1x2x3
(b) Topology abstraction.
Figure 2: A typical T-intersection. Six connecting roads link three 2-lane roads in the junction area.
4 Traffic Scenario Logic for Highway
@@ 4.1 Abstract Model
This section formally defines a scenario on a uni-direction multi-lane highway, where one or more vehicles travel on a road with multiple lanes. The road length is infinite, and the number of lanes remains constant throughout. Vehicles can perform lane-changing maneuvers between lanes to switch to adjacent lanes.
@@
The lateral position of a vehicle (perpendicular to the direction of the road) is described using the lanes it occupies, while the longitudinal position (along the direction of the road) can be described using the positional relationship with other vehicles. Using the road coordinate system, it can be defined as follows: {definition}[Longitudinal Positional Relationship] Denote the range occupied by vehicle $c$ as $[s_{c}^{-},s_{c}^{+}]$, where $s_{c}^{-}$ and $s_{c}^{+}$ represent the $s$-coordinates of the rear and front ends of the vehicle, respectively. Assuming that the $S$ axis direction aligns with the direction of the vehicle, based on the road coordinate system, we can categorize the longitudinal positional relationship between two vehicles $c_{1}$ and $c_{2}$ on the same road as follows:
@@
    Cover: If $s_{c_{1}}^{-}\leq s_{c_{2}}^{+}$ and $s_{c_{1}}^{+}\geq s_{c_{2}}^{-}$, then $c_{1}$ is covering $c_{2}$.
@@
The definition is similar when the $S$ axis direction is opposite the direction of the vehicle.
@@
For convenience, further define the tail of a scenario as: {definition}[Tail of a Scenario] The tail of scenario $\mathcal{S}$, denoted by $\mathcal{S}_{i:}=(\mathbb{C},\mathbb{L},\mathbb{D},T-i,\langle\mathfrak{s}_{i},\mathfrak{s}_{i+1},\dots,\mathfrak{s}_{T-1}\rangle)$ $(0\leq i<T)$
@@
where, $c,c_{1},c_{2}\in\mathbb{C}$, $l,l_{1},l_{2}\in\mathbb{L}$, $d,d_{1},d_{2}\in\mathbb{D}$. The modal operators and take their meanings “Next” and “Globally” in temporal logic, respectively. $\mathbfbb{F}$ denotes that the length of this scenario is 1, indicating the final state. {remark} We can further define $\phi\land\psi$ , $\phi\lor\psi$ , $\exists x\,\phi$ and $\mdwhtlozenge\phi$ as ($(\lnot(\phi\to(\lnot\psi)))$ , ((¬$((\lnot\phi)\to\psi)$ , ($(\lnot\forall x\,(\lnot\phi))$ and ($(\lnot\mdwhtsquare\lnot\phi)$.
@@
[Semantics of Highway TSL] Given a road $\mathcal{R}$ and a scenario $\mathcal{S}$, the satisfaction of formulae in the subsection 4.1 follows the semantics listed below:
@@ 4.2 Rules of Highway TSL
Realistic scenarios follow physics rules and traffic rules. In this subsection, we list these rules in the form of TSL formulae, where the traffic rules can be added or subtracted according to the laws of different regions.
Physics Rules
  1. 1.
    The longitudinal positional relationships are symmetric
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\to\mathsf{lonr}(c_{2},c_{1},\mathsf{behind})),$
@@
  2. 2.
@@
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall c_{3}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\\ \land\mathsf{lonr}(c_{2},c_{3},\mathsf{ahead})\to\mathsf{lonr}(c_{1},c_{3},\mathsf{ahead})),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall c_{3}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\\ \land\mathsf{lonr}(c_{2},c_{3},\mathsf{behind})\to\mathsf{lonr}(c_{1},c_{3},\mathsf{behind})).$
  3. 3.
    Contradictory longitudinal positional relationships.
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall c_{3}\,\lnot(\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\\ \land\mathsf{lonr}(c_{2},c_{3},\mathsf{cover})\land\mathsf{lonr}(c_{3},c_{1},\mathsf{ahead}))),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall c_{3}\,\lnot(\mathsf{lonr}(c_{1},c_{2},\mathsf{cover})\\ \land\mathsf{lonr}(c_{2},c_{3},\mathsf{ahead})\land\mathsf{lonr}(c_{3},c_{1},\mathsf{ahead}))).$
  4. 4.

### v3

@@ Abstract
Formal representations of traffic scenarios can be used to generate test cases for the safety verification of autonomous driving. However, most existing methods are limited to highway or highly simplified intersection scenarios due to the intricacy and diversity of traffic scenarios. In response, we propose Traffic Scenario Logic (TSL), which is a spatial-temporal logic designed for modeling and reasoning of urban pedestrian-free traffic scenarios. TSL provides a formal representation of the urban road network that can be derived from OpenDRIVE, i.e., the de facto industry standard of high-definition maps for autonomous driving, enabling the representation of a broad range of traffic scenarios without discretization approximations. We implemented the reasoning of TSL using Telingo, i.e., a solver for temporal programs based on Answer Set Programming, and tested it on different urban road layouts. Demonstrations show the effectiveness of TSL in test scenario generation and its potential value in areas like decision-making and control verification of autonomous driving. The code for TSL reasoning has been open-sourced.
Codehttps://github.com/USTCAVSA/TSL
@@ 1 Introduction
Autonomous driving technology is promising to revolutionize transportation by enhancing safety, efficiency, and accessibility. The rigorous testing and verifying their safety under various traffic scenarios is central to developing and deploying autonomous vehicles. Formal representations of these scenarios are crucial for generating comprehensive test cases and ensuring the reliability of autonomous driving systems (Zhang et al. 2021).
Existing methods for formalizing traffic scenarios often face limitations, particularly in modeling intricate urban environments with diverse traffic patterns. While some approaches excel in representing highway scenarios or simplified intersections, they struggle to capture the complexity inherent in urban driving environments (Hilscher and Schwammberger 2016; Yao and Li 2023; Li, Tao, and Wotawa 2020). Consequently, there is a pressing need for a versatile and robust formal logic framework capable of modeling a broad spectrum of urban traffic scenarios.
In response to this challenge, we introduce Traffic Scenario Logic (TSL), a novel spatial-temporal logic specifically tailored for modeling and reasoning about urban pedestrian-free traffic scenarios. Unlike previous methods, TSL is designed to provide a comprehensive formal representation of the urban road network without discretization approximations, enabling the modeling of various traffic configurations encountered in complex urban environments.
One of the main benefits of TSL is its compatibility with OpenDRIVE, which is the industry-standard format used for high-definition (HD) maps in autonomous driving (Dupuis, Strobl, and Grezlikowski 2010). By utilizing OpenDRIVE data, TSL seamlessly translates real-world road networks into logical representations, enabling the representation of a wide range of urban traffic scenarios.
TSL also distinguishes itself with a unique spatial representation. Instead of using a simple discretized position description, TSL captures a vehicle’s longitudinal position based on its relationship with other vehicles. By merging scenarios with the same vehicle relationships, TSL improves efficiency without losing comprehensive coverage of scenarios.
We first give the syntax and semantics of TSL for uni-direction multi-lane highway scenarios in Section 3, then extend it to urban scenarios in Section 4. Finally, we implement the reasoning of TSL with Telingo (Cabalar et al. 2019), a solver for temporal programs based on Answer Set Programming. Section 5 demonstrates several examples on different urban road layouts, showing the effectiveness of TSL in test scenario generation and potential applications in decision-making and control verification of autonomous driving.
@@
  • Provides a formal representation of an urban road network that is compatible with the OpenDRIVE HD map format and an efficient and complete spatial representation.
  • Implements the reasoning mechanism with Telingo solver and provides examples of different road layouts.
@@ 2 Related Work
Hilscher et al. (2011) were one of the first few work on using spatial-temporal logic to model traffic scenarios, proposing Multi-Lane Spatial Logic (MLSL) which provides an abstract model of a uni-directional multi-lane motorway. In MLSL, vehicles’ position, speed, and acceleration constitute a traffic snapshot, and the traffic snapshots at different moments constitute a traffic scenario. Changes between traffic snapshots are described by transition rules, inspired by $\mathcal{ITL}$ (Moszkowski 1982). Hilscher, Linker, and Olderog (2013) proposed an extended version of MLSL that supports a bi-directional motorway and length measurement for road segments. Hilscher and Schwammberger (2016) further proposed Urban Multi-Lane Spatial Logic (UMLSL) which contains the modeling of intersections, simplifying an intersection of two two-way, single-lane highways into four crossing road segments. Meanwhile, Xu and Li (2016) introduced vertical lanes that intersect with the horizontal lanes by extending the position to two dimensions to create an urban road network, formalized in a grid way. Based on UMLSL, Schwammberger and Alves (2021) proposed Urban Spatial Logic for Traffic Rules (USL-TR) to add static objects and non-autonomous road users. Xu et al. (2019) represented an urban road network with road segments and the arcs that link them together, which was further formalized into a hierarchical structure by Yao and Li (2023). Some other research includes modeling an intersection controlled by traffic lights (Loos and Platzer 2011); using the State-Clock Logic for the representation of time (Bischopink and Olderog 2022); and employing ontology (Bagschik, Menzel, and Maurer 2018; Li, Tao, and Wotawa 2020).
@@
The above methods for describing spatial relationships rely on either road segments or a grid-based occupancy approach, which have their limitations. Distinguishing between consecutive free grids or road segments is usually not necessary, and it can be crucial to describe the positional relationships of vehicles within the same grid or road segment. In this regard, a simple discretized description of the road network is often insufficient for test scenario generation.
3 Highway Traffic Scenario Logic
@@ 3.1 Abstract Model
This section formally defines a scenario on a uni-direction multi-lane highway, where one or more vehicles travel on a road with multiple lanes. The road length is infinite, and the number of lanes remains constant throughout. Please note that a road with variable lane numbers can be modeled as separate roads, which will be discussed in Section 4. Vehicles can perform lane-changing maneuvers between lanes to switch to adjacent lanes.
@@
The lateral position of a vehicle (perpendicular to the direction of the road) is described using the lanes it occupies, while the longitudinal position (along the direction of the road) can be described using the positional relationship with other vehicles. Using the road coordinate system, it can be defined as follows: {definition}[Longitudinal Positional Relationship] Denote the range occupied by vehicle $c$ as $[s_{c}^{-},s_{c}^{+}]$, where $s_{c}^{-}$ and $s_{c}^{+}$ represent the $s$-coordinates of the rear and front ends of the vehicle under Frenet coordinate system (Werling et al. 2010), respectively. Assuming that the $S$ axis direction aligns with the direction of the vehicle, we can categorize the longitudinal positional relationship between two vehicles $c_{1}$ and $c_{2}$ on the same road as follows:
@@
    Cover: If $s_{c_{1}}^{-}\leq s_{c_{2}}^{+}$ and $s_{c_{1}}^{+}\geq s_{c_{2}}^{-}$, then $c_{1}$ covers $c_{2}$.
@@
The definition is similar when the $S$ axis direction is opposite to that of the vehicles. Obviously, their longitudinal positional relationship falls into one of these three categories for two vehicles on the same road.
@@
For convenience, further define the $i$-tail of a scenario as: {definition}[$i$-tail of a Scenario] The $i$-tail of scenario $\mathcal{S}$, denoted by $\mathcal{S}_{i:}$, is defined as $(\mathbb{C},\mathbb{L},\mathbb{D},T-i,\langle\mathfrak{s}_{i},\mathfrak{s}_{i+1},\dots,\mathfrak{s}_{T-1}\rangle)$ for all $0\leq i<T$
@@
where, $c,c_{1},c_{2}\in\mathbb{C}$, $l,l_{1},l_{2}\in\mathbb{L}$, $d,d_{1},d_{2}\in\mathbb{D}$. The modal operators and take their meanings “Next” and “Globally” in temporal logic, respectively. $\mathbfbb{F}$ denotes that the length of this scenario is 1, indicating the final state. {remark} We can further define $\phi\land\psi$ , $\phi\lor\psi$ , $\exists x\,\phi$ and $\mdwhtlozenge\phi$ as $\lnot(\phi\to(\lnot\psi))$ , ($(\lnot\phi)\to\psi$ , $\lnot\forall x\,(\lnot\phi)$ and $\lnot\mdwhtsquare\lnot\phi$, respectively.
@@
[Semantics of Highway TSL] Given a road $\mathcal{R}$ and a scenario $\mathcal{S}$, the satisfaction of formulae in the Section 3.1 follows the semantics listed below:
@@ 3.2 Rules of Highway TSL
Realistic scenarios obey both physics rules and traffic rules. In this subsection, we outline these rules and auxiliary predicates we use in the form of TSL formulae, as shown in Table 1. Traffic rules can be added or subtracted according to the laws of different regions.
Rule TSL Formula
PR1 $\mdwhtsquare(\forall c_{1}\forall c_{2}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\to\mathsf{lonr}(c_{2},c_{1},\mathsf{behind}))$,
@@
PR2 $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall c_{3}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\land\mathsf{lonr}(c_{2},c_{3},\mathsf{ahead})\to\mathsf{lonr}(c_{1},c_{3},\mathsf{ahead}))$,
$\mdwhtsquare(\forall c_{1}\forall c_{2}\forall c_{3}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\land\mathsf{lonr}(c_{2},c_{3},\mathsf{behind})\to\mathsf{lonr}(c_{1},c_{3},\mathsf{behind}))$.
PR3 $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall c_{3}\,\lnot(\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\land\mathsf{lonr}(c_{2},c_{3},\mathsf{cover})\land\mathsf{lonr}(c_{3},c_{1},\mathsf{ahead})))$.
PR4 $\mathbfbb{F}\lor\mdwhtsquare(\forall c_{1}\forall c_{2}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\to\mdwhtcircle(\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\lor\mathsf{lonr}(c_{1},c_{2},\mathsf{cover})))$,
$\mathbfbb{F}\lor\mdwhtsquare(\forall c_{1}\forall c_{2}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\to\mdwhtcircle(\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\lor\mathsf{lonr}(c_{1},c_{2},\mathsf{cover})))$.
PR5 $\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\forall l_{3}\,\mathsf{on}(c,l_{1})\land\mathsf{on}(c,l_{2})\land\mathsf{cleft}(l_{1},l_{3})\land\mathsf{cleft}(l_{3},l_{2})\to\mathsf{on}(c,l_{3}))$.
PR6 $\mdwhtsquare(\forall c\exists l\,\mathsf{on}(c,l))$.
PR7 $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\,\lnot(\mathsf{on}(c,l_{1})\land\mathsf{on}(c,l_{2})\land\lnot(l_{1}=l_{2})\land\lnot\mdwhtcircle\mathsf{on}(c,l_{1})\land\lnot\mdwhtcircle\mathsf{on}(c,l_{2})))$,
$\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\,\lnot(\lnot\mathsf{on}(c,l_{1})\land\lnot\mathsf{on}(c,l_{2})\land\lnot(l_{1}=l_{2})\land\mdwhtcircle\mathsf{on}(c,l_{1})\land\mdwhtcircle\mathsf{on}(c,l_{2})))$,
$\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\,\lnot(\mathsf{on}(c,l_{1})\land\lnot\mathsf{on}(c,l_{2})\land\lnot(l_{1}=l_{2})\land\lnot\mdwhtcircle\mathsf{on}(c,l_{1})\land\mdwhtcircle\mathsf{on}(c,l_{2})))$.
TR1 $\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\forall l_{3}\,\lnot(\mathsf{on}(c,l_{1})\land\mathsf{on}(c,l_{2})\land\mathsf{on}(c,l_{3})\land\lnot(l_{1}=l_{2})\land\lnot(l_{1}=l_{3})\land\lnot(l_{2}=l_{3})))$.
TR2 $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall l\,\lnot(\mathsf{on}(c_{1},l)\land\mathsf{on}(c_{2},l)\land\mathsf{lonr}(c_{1},c_{2},\mathsf{cover})\land\lnot(c_{1}=c_{2})))$.
AP1 $\forall l_{1}\forall l_{2}\,\mathsf{left}(l_{1},l_{2})\to\mathsf{cleft}(l_{1},l_{2})$,
$\forall l_{1}\forall l_{2}\forall l_{3}\,\mathsf{left}(l_{2},l_{3})\land\mathsf{cleft}(l_{1},l_{2})\to\mathsf{cleft}(l_{1},l_{3})$.
Table 1: The rules and auxiliary predicates used in Highway TSL
Physics Rules
  1. PR1
    The longitudinal positional relationships are symmetric.
  2. PR2
@@
  3. PR3
    Longitudinal positional relationships must not be contradictory.
  4. PR4